So I would think that the average of4,2,-3,6,5,-3,-2,0,-1,-4,3 would be .636363 but the tester said it is 1.18 no ClUE! I came up with an answer of 1.18 by adding only the negative numbers and dividing by the whole total of numbers! but WHY! and is that the way it is supposed to be done.

Answer by checkley71(8403)

4+2+6+5+0+3=20 & -3-3-2-1-4=-13 or 20-13=7 the total value from the 11 numbers is 7 thus the average is 7/11 or .636363
this the correct answer if these are the 11 numbers and you were asked to find the average of all numbers.
however if you were asked to find the average of the negative numbers that would be -13/11 or -1.181818